# How Do Developers Interact with AI? An Exploratory Study on Modeling Developer Programming Behavior

### 28 Mar 2026 v1

@@ Abstract.
Artificial Intelligence (AI) is reshaping how developers adopt software engineering practices, yet the multi-dimensional nature of developer-AI interaction remains under-explored. Prior studies have primarily examined dimensions observable from developer activities such as “Prompt crafting”, and “Code Editing”, overlooking how hidden intentions and emotional dimensions intertwine with concrete actions during AI-assisted programming. Understanding the interplay is essential for improving developer experience and future AI assistant designs. To understand this phenomenon, we conducted a mixed-methods study with 76 developers. We first split developers into AI-assisted and non-AI groups. Each performed programming tasks (Python with API management or Java with SQL). Developers retrospectively labeled their self-reported intentions, tool-supported actions, and emotions (on a 7-point valence scale) from screen recordings, supplemented by participant surveys and interviews. Our user study resulted in a novel model named S-IASE with four dimensions to describe a programming behavior: intention, action, supporting tool, and emotion for a given development state. Our analysis reveals several aggregated and sequential behavioral patterns. For example, for aggregated patterns, using AI assistants often makes developers more focused on actively “creating” code, evaluating, and verifying these generated results; for sequential pattern, AI-assisted participants showed emotionally stable development flow, as opposed to non-AI-assisted participants who experienced more fluctuated emotions. Interviews revealed further nuance: some developers reported impostor-like feelings, expressing guilt or self-doubt about relying on AI for programming. The uncovered patterns indicate our model can provide actionable insights for improving AI assistants’ responsiveness, training developers in AI collaboration, and designing developer-centric AI studies. Our work bridges an important gap in understanding the complexities of developer-AI interaction in programming context and also sheds light on future developer centric research directions.
copyright: none
@@ 1. Introduction
Software engineering is undergoing a significant shift due to the rise of AI assistants (Kam et al., 2025; Zakharov et al., 2025; Chen et al., 2021). AI assistants such as GitHub Copilot (Ray, 2023) and ChatGPT (Roth, 2024) have gained widespread attention and quickly gained millions of subscribers from students alike to professional developers. GitHub Copilot, in particular, promises to be an “AI [code] editor for everyone” (22). Prior research has explored these AI assistants in various capacities to support software developers throughout the software development lifecycle, ranging from requirements (Terragni et al., 2025; Robinson et al., 2024), design (Robles-Aguilar et al., 2021), development (Torka and Albayrak, 2024; Carleton et al., 2024; Yetiştiren et al., 2023; Murali et al., 2024), testing (Aleti, 2023), and even maintenance (Marchezan et al., 2024; Bidollahkhani and Kunkel, 2024). Due to the wide variety of areas in which a developer could leverage AI, recent large-scale developer surveys found that 76% of developers in 2024 (2) and 84% of developers in 2025 are already using or plan to use AI in their work (1).
Yet, as AI assistants become embedded in everyday software programming, a central question arises: how do developers interact with these assistants? Without sufficient understanding, we cannot design AI assistants that better align with developer programming behavior and ultimately boost developers’ productivity.
@@
Most prior research has focused on ostensible benefits (e.g., reduced repetitive coding  (Ng et al., 2024)) or on narrow contexts (e.g., novice learning (Amoozadeh et al., 2024; Kazemitabaar et al., 2023)). Even when prior studies have looked at interaction, they have often done so from the researchers’ observation, identifying a limited set of modes (e.g., “acceleration” vs. “exploration” modes (Barke et al., 2023)). However, few studies have focused on developer-AI interaction from the developer’s perspective. We lack perspective on what developers think and feel when they interact with AI assistants as well as how those perceptions differ from those without AI.
@@
Mozannar et al. (Mozannar et al., 2024) are among the first to explore developer-AI interaction from the developer’s perspective. They investigated how developers interact with GitHub Copilot during time-restricted coding tasks and developed a taxonomy called CUPS to categorize developer states when interacting with AI assistants. While their work represents a significant step toward understanding developer-AI interaction, the CUPS taxonomy has treated a “state” as a single coding activity and focused on transitions between activities (e.g., from “Prompt crafting” to “Code Editing”).
However, such activity-based labels failed to capture the subtleties and multidimensionality of developer-AI interaction in the context of programming.
@@
By “dimension”, we refer to a distinct, analyzable component of developer-AI interaction (e.g., intentions, actions). Each dimension provides a lens to study and compare interactions. Google’s SIA (State–Intention–Action) model (Research, 2023) further inspired us to view a programming state not as a single coding activity, but as a snapshot of development context, which includes the specific tools, and that gives rise to intentions and actions. These insights led to our understanding that developer-AI programming behavior involves more than one dimension.
This led to the main question guiding our research:
@@
Our exploratory work aims to understand more deeply how AI assistants shape the different dimensions of developer programming behavior and how these dimensions could be characterized to the sequences of development states.
We conducted a mixed-methods user study with a group of 76 participants with varied background in terms of education, programming experience, and familiarities with the topic of the tasks. The participants were assigned to two groups: AI and non-AI. Participants from the AI group could freely use any AI assistants in their preferred way during task execution. In contrast, those from the non-AI group, who had little to no experience with using AI for programming in the past, performed the task without using AI. Within each group, participants were assigned to one of two programming tasks: a Java task with MySQL Database Operations or a Python task with GitHub REST API.
From this study, we found that developer programming behavior when interacting with AI involves four distinct dimensions.
We formalize this in our proposed S-IASE model, which characterizes a developer’s State when they interact with AI assistants, using their Intention, pairs of ¡Action, Support tool¿, and Emotion.
@@
Our model details how intention (e.g., To Implement Code $\to$ To Resolve Errors), actions supported with tools (e.g., ¡Reading and Comprehension, No Tool¿ $\to$ ¡Crafting Query/Prompt, AI Tool¿), and emotions (e.g., Neutral $\to$ Positive) manifest throughout developers’ states.
Our proposed model unlocks the new opportunity to perform analysis and reveal behavioral patterns. Specifically, on the top of S-IASE, we performed mixed-methods analysis and uncover two complementary types of patterns. First, aggregated patterns derived from descriptive statistics. For instance, our results show that using AI assistants often makes the programming experience more focused on actively “creating” code, evaluating, and verifying these generated results. Second, sequential patterns identified through automated sequential pattern mining approaches (e.g., CloSpan (Yan et al., 2003)). The identified patterns indicate that AI-assisted participants exhibited more emotionally stable development flows, whereas non-AI participants experienced more fluctuating emotions. Finally, our interview thematic analysis based on socio-technical grounded theory (STGT) (Hoda, 2024) revealed more subtle findings. Despite the statistical stability, some AI-assisted participants reported impostor-like feelings, expressing guilt or self-doubt about relying on AI. These qualitative findings complement the quantitative results, highlighting the complex emotional trade-offs in AI-assisted development.
Our work serves as an important exploration of the deeper complexities of developer-AI interaction in software development. The carefully constructed model and insights drawn from our user study provide valuable guidance for both researchers and practitioners, informing the design of future studies and the development of more effective tools for AI-assisted programming.
The main contributions of our work include:
  • A model named S-IASE for describing programming behavior which covers four dimensions: intention, action, supporting tools, and emotion to a given sequential state when developing software.
@@
  • Analysis to uncover both aggregated and sequential programming behavior patterns based on our S-IASE model.
@@ 2.1. Research Questions
RQ2: What aggregated behavioral patterns emerge for developers-AI interactions?
To answer this, we performed statistical analysis on the distribution of intentions, actions, tools, and emotions across AI-assisted and non-AI groups, identifying consistent differences across tasks.
RQ3: What sequential behavioral patterns emerge for developers-AI interactions?
We applied the closed sequential pattern mining algorithm (CloSpan) to annotated sequences, uncovering recurring sequences for each dimension. Then we conducted a thematic analysis using socio-technical grounded theory (STGT) on interviews to capture developers’ perceptions and subtle experiences with AI assistance.
@@ 2.2. S-IASE Model Construction
Stage 1: Initialization with observations.
We built a structured model with multiple distinct dimensions, with each dimension representing a perspective along which interaction can be studied and compared. Our design was inspired by a State–Intent–Action (SIA) model proposed by a Google Research team to characterize software development activities (Research, 2023). It includes but is not limited to three broad categories of developers’ activities: (1) debugging and repair tasks, (2) code review tasks, and (3) code editing tasks. However, the original SIA model is not open-sourced and does not disclose the important details, such as category definitions and whether a user study was conducted internally at Google.
Moreover, the SIA model is code-centric by design. Differently, our study must capture concrete, context-rich behavioral patterns in the programming context from a human-centric perspective to draw conclusions about developer actions and emotions.
Stage 2: Model Adaptation. We developed a new model with additional dimensions and provide a clearer definition of each dimension. The transition from Google’s SIA model to our proposed S-IASE model was driven by both empirical evidence from our pilot study and established literature. One new dimension is paired with action, named supporting tools. This dimension was introduced based on the insight that developers in a production environment do not rely on AI exclusively; rather, they conduct actions by navigating a heterogeneous tool ecosystem. We therefore use supporting tools to distinguish between actions done with or without AI, or any other tools. We also added the emotions dimension, informed by prior literature (Cheng et al., 2023; Girardi et al., 2020; Graziotin et al., 2015; Madampe et al., 2025) and our pilot study (Section 2.7) to account for affective states that can reflect how emotions and cognitive programming behavior interplay (Girardi et al., 2020) and shape together the multidimensionality of developer-AI interactions. This results in our S–IASE model (State, Intention, Action, Supporting Tools, Emotion).
Stage 3: Dimension labeling. To develop the labels within each dimension, we drafted an initial set of labels based on prior studies (Mozannar et al., 2024; Amoozadeh et al., 2024; Russell, 1980; Gardella et al., 2024). Then, we allowed our participants to add their custom labels during annotation if the predefined options were insufficient.
@@ 2.3. User Study Procedure
S1: Pre-survey. Each participant completed a pre-survey to collect their consent to participate, demographic information (e.g., educational background), programming experience (especially on their AI use), and their prior knowledge on the overall topic of our designed tasks (e.g., familiarity with writing and executing SQL queries). We then used the information as a guide to assign participants to different groups (described in Section 2.5).
@@
S2: Activity. We provided two options for Setup according to participants’ preferences: (1) installing essential tools needed for the activity on their computers locally, such as task-specific programming language interpreters, a necessary database, a screen recording software, and our activity annotation tool. (2) installing a virtual machine that contains all these essential tools. Participants assigned to the AI-assisted group were explicitly informed that they could freely install and use their preferred AI assistants (e.g., ChatGPT or GitHub Copilot) without restriction, whereas participants in the non-AI group were instructed not to use any AI assistants during the tasks. Then, we introduced participants to the tasks and group assignment (i.e., AI or non-AI). Participants started screen recording and then performed their assigned task (Task Execution). After the tasks were completed, each participant watched their screen recording and used our custom annotation tool to break down the whole video into a sequence of intervals. For each interval, participants were asked to label their intentions, actions, supporting tools, and emotions.
S3: Post-survey. We first asked participants to upload their (1) screen recording of their task execution, (2) annotation results, and (3) implemented code. Then, we asked participants to reflect on their experience in completing the task in terms of cognitive load, self-efficacy, and challenges during task execution. Lastly, we collected their willingness to participate in the follow-up interview.
S4: Interview (optional). We invited interested participants to schedule an interview. These helped triangulate our findings from the prior three stages. This additional data collection helped to develop a deeper understanding of participant behavior.
@@ 2.4. Activity Tasks
$\bullet$ JavaT: Java Task with MySQL Database Operations This task involves setting up their MySQL database, implementing Java database operations to interact with the MySQL database, and solving several business queries over 4 subtasks.
@@ 2.5. Participants and Task Assignment
We recruited participants for our user study following in two ways. First was an in-class mode, where we gave a presentation in both undergraduate- and graduate-level software engineering courses to explain the study’s objectives and procedures. Second was recruiting participants through the research team’s internal connections and snowball sampling. To become a participant, one must be over 18 years old and have prior programming experience. Our recruitment resulted in 90 participants, with 11 participants recruited via snowballing and 79 from the classroom. As compensation, 8 randomly selected participants received \$25 Amazon gift cards.
@@
We used stratified sampling to assign each participant to a group by following the criteria: (1) If a participant has never or rarely used AI for programming, we assigned them to non-AI groups, i.e., nonAI-PyT and nonAI-JavaT. (2) Following the practices from Mozanner et al. (Mozannar et al., 2024), we assigned participants to a task according to their prior task-related knowledge in the pre-survey. We therefore created four participant groups: AI-PyT: participants in this group solved PyT and were allowed to use AI, nonAI-PyT: participants in this group solved PyT and were not allowed to use AI, AI-JavaT: participants in this group solved JavaT and were allowed to use AI, nonAI-JavaT: participants in this group solved JavaT and were not allowed to use AI.
For data filtering, we do not consider a participant’s data if the screen recording did not capture the full screen or the annotation log file went missing. Moreover, to prevent the use of secondary devices (e.g., participants in non-AI groups using personal phones to access AI), researchers were present during all sessions. We also excluded four participants who did not complete the tasks during class. Researchers thoroughly examined the videos that participants recorded, and we excluded one participant in the non-AI groups who used AI in experiments. After filtering, we were left with 76 participants who completed all the study steps and complied with our instructions. Over 76% of participants reported 3 or more years of programming experience, with a median of over 4 years, and 75% reported prior internship or industry experience. Full demographic information is available in our replication package (Wu et al., 2025).

### 26 Apr 2026 v2

@@ Abstract.
Artificial Intelligence (AI) is reshaping how developers adopt software engineering practices, yet the multi-dimensional nature of developer-AI interaction remains under-explored. Prior studies have primarily examined dimensions observable from developer activities such as “Prompt Crafting” and “Code Editing,” overlooking how hidden intentions and emotional dimensions intertwine with concrete actions during AI-assisted programming. Understanding the interplay is essential for improving developer experience and future AI assistant designs. To understand this phenomenon, we conducted a mixed-methods study with 76 developers. We first split developers into AI-assisted and non-AI groups. Each developer performed a programming task (either Python with API management or Java with SQL). Developers retrospectively labeled their self-reported intentions, tool-supported actions, and emotions (on a 7-point valence scale) from screen recordings, supplemented by participant surveys and interviews. Our user study resulted in a novel model, named S-IASE, with four dimensions to describe programming behavior for a given development state: intention, action, supporting tool, and emotion. Our analysis reveals several aggregated and sequential behavioral patterns. For example, for aggregated patterns, using AI assistants often led developers to focus more on actively “creating” code, evaluating, and verifying the generated results; for sequential patterns, AI-assisted participants showed emotionally stable development flows, as opposed to non-AI-assisted participants who experienced more fluctuating emotions. Interviews revealed further nuance: some developers reported impostor-like feelings, expressing guilt or self-doubt about relying on AI for programming. The uncovered patterns indicate that our model can provide actionable insights for improving AI assistants’ responsiveness, training developers in AI collaboration, and designing developer-centric AI studies. Our work bridges an important gap in understanding the complexities of developer-AI interaction in the programming context and sheds light on future developer-centric research directions.
AI-assisted programming, programming behavior, human-AI collaboration, developer experience
copyright: ccdoi: 10.1145/3808120journalyear: 2026journal: PACMSEjournalvolume: 3journalnumber: FSEarticle: FSE113publicationmonth: 7submissionid: fse26mainb-p272-pccs: Software and its engineering Collaboration in software developmentccs: Human-centered computing Human computer interaction (HCI)$c$s: Computing methodologies Artificial intelligence
@@ 1. Introduction
Software engineering is undergoing a significant shift due to the rise of AI assistants (Kam et al., 2025; Zakharov et al., 2025; Chen et al., 2021). AI assistants such as GitHub Copilot (Ray, 2023) and ChatGPT (Roth, 2024) have gained widespread attention and quickly attracted millions of subscribers from students to professional developers alike. GitHub Copilot, in particular, promises to be an “AI [code] editor for everyone” (22). Prior research has explored these AI assistants in various capacities to support software developers throughout the software development lifecycle, ranging from requirements (Terragni et al., 2025; Robinson et al., 2025), design (Robles-Aguilar et al., 2021), development (Torka and Albayrak, 2024; Carleton et al., 2024; Yetiştiren et al., 2023; Murali et al., 2024), testing (Aleti, 2023), to even maintenance (Marchezan et al., 2024; Bidollahkhani and Kunkel, 2024). Due to the wide variety of areas in which a developer can leverage AI, recent large-scale developer surveys found that 76% of developers in 2024 (2) and 84% of developers in 2025 are already using or plan to use AI in their work (1).
@@
Yet, as AI assistants become embedded in everyday software programming, a central question arises: how do developers interact with these assistants? Without sufficient understanding, we cannot design AI assistants that better align with developer programming behavior and ultimately boost developers’ productivity. Most prior research has focused on ostensible benefits (e.g., reduced repetitive coding  (Ng et al., 2024)) or on narrow contexts (e.g., novice learning (Amoozadeh et al., 2024; Kazemitabaar et al., 2023)). Even when prior studies have looked at interaction, they have often done so from the researchers’ observations, identifying a limited set of modes (e.g., “acceleration” vs. “exploration” modes (Barke et al., 2023)). However, few studies have focused on developer-AI interaction from the developer’s perspective. We have limited perspectives on what developers think and feel when they interact with AI assistants as well as how those perceptions differ from those without AI.
@@
Mozannar et al. (Mozannar et al., 2024) are among the first to explore developer-AI interaction from the developer’s perspective. They investigated how developers interact with GitHub Copilot during time-restricted coding tasks and developed a taxonomy called CUPS to categorize developer states when interacting with AI assistants. While their work represents a significant step toward understanding developer-AI interaction, the CUPS taxonomy treated a “state” as a single coding activity and focused on transitions between activities (e.g., from “Prompt Crafting” to “Code Editing”). However, such activity-based labels failed to capture the subtleties and multidimensionality of developer-AI interaction in the context of programming.
@@
By “dimension,” we refer to a distinct, analyzable component of developer-AI interaction (e.g., intentions, actions). Each dimension provides a lens to study and compare interactions. Google’s SIA (State–Intention–Action) model (Research, 2023) further inspired us to view a programming state not as a single coding activity, but as a snapshot of development context, which includes the specific tools, and gives rise to intentions and actions. These insights led to our understanding that developer-AI programming behavior involves more than one dimension.
This motivated the main question guiding our research:
@@
Our exploratory work aims to understand more deeply how AI assistants shape the different dimensions of developer programming behavior and how these dimensions could be characterized to the sequences of development states. We conducted a mixed-methods user study with a group of 76 participants with varied backgrounds in terms of education, programming experience, and familiarity with the topic of the tasks. The participants were assigned to two groups: AI and non-AI. Participants from the AI group could freely use any AI assistants in their preferred way during task execution. In contrast, those from the non-AI group, who had little to no experience with using AI for programming in the past, performed the task without using AI. Within each group, participants were assigned to one of two programming tasks: a Java task with MySQL Database Operations or a Python task with GitHub REST API.
From this study, we found that developer programming behavior when interacting with AI involves four distinct dimensions. We formalize this in our proposed S-IASE model, which characterizes a developer’s State when they interact with AI assistants, using their Intention, pairs of ¡Action, Supporting tool¿, and Emotion.
@@
Our model details how intention (e.g., To Implement Code $\to$ To Resolve Errors), actions supported with tools (e.g., ¡Reading and Comprehension, No Tool¿ $\to$ ¡Crafting Query/Prompt, AI Tool¿), and emotions (e.g., Neutral $\to$ Positive) manifest throughout developers’ states. Our proposed model unlocks the new opportunity to perform analysis and reveal behavioral patterns. Specifically, on top of S-IASE, we performed a mixed-methods analysis and uncovered two complementary types of patterns. First, aggregated patterns derived from descriptive statistics. For instance, our results show that using AI assistants often makes the programming experience more focused on actively “creating” code, evaluating, and verifying the generated results. Second, sequential patterns identified through automated sequential pattern mining approaches (e.g., CloSpan (Yan et al., 2003)). The identified patterns indicate that AI-assisted participants exhibited statistically more stable emotional development flows, whereas non-AI participants experienced more fluctuating emotions. Finally, our interview thematic analysis based on socio-technical grounded theory (STGT) (Hoda, 2024) revealed more subtle findings: despite this statistical stability, some AI-assisted participants reported impostor-like feelings, expressing guilt or self-doubt about relying on AI. These findings highlight the complex emotional trade-offs in AI-assisted development.
Our work explores the deeper complexities of developer-AI interaction in software development. The carefully constructed model and insights drawn from our user study provide valuable guidance for both researchers and practitioners, informing the design of future studies and the development of more effective tools for AI-assisted programming. The main contributions of our work include:
  • A model named S-IASE for describing programming behavior, which covers four dimensions: intention, action, supporting tool, and emotion to a given sequential state.
@@
  • An Analysis to uncover both aggregated and sequential programming behavior patterns based on our S-IASE model.
@@ 2.1. Research Questions
RQ2: What aggregated behavioral patterns emerge for developer-AI interactions?
To answer this, we performed a statistical analysis on the distribution of intentions, actions, tools, and emotions across AI-assisted and non-AI groups, identifying consistent differences across tasks.
RQ3: What sequential behavioral patterns emerge for developer-AI interactions?
We applied the closed sequential pattern mining algorithm (CloSpan) to the annotated sequences, uncovering recurring sequences for each dimension. Then we conducted a thematic analysis using STGT (Hoda, 2024) on the interviews to capture developers’ perceptions and experiences with AI assistance.
@@ 2.2. S-IASE Model Construction
Stage 1: Initialization with observations. We built a structured model with multiple distinct dimensions, with each dimension representing a perspective along which interaction can be studied and compared. Our design was inspired by a State–Intent–Action (SIA) model proposed by a Google Research team to characterize software development activities (Research, 2023). It includes but is not limited to three broad categories of developers’ activities: (1) debugging and repair tasks, (2) code review tasks, and (3) code editing tasks. However, the original SIA model is not open-sourced and does not disclose the important details, such as category definitions and whether a user study was conducted internally at Google. Moreover, the SIA model is code-centric by design. Differently, our study must capture concrete, context-rich behavioral patterns in the programming context from a human-centric perspective to draw conclusions about developer actions and emotions.
Stage 2: Model Adaptation. We developed a new model with additional dimensions and provided clearer definitions of each dimension. The transition from Google’s SIA model to our proposed S-IASE model was driven by both empirical evidence from our pilot study and established literature. One new dimension, named supporting tools, is paired with action. This dimension was introduced based on the insight that developers in a production environment do not rely on AI exclusively; rather, they conduct actions by navigating a heterogeneous tool ecosystem. We therefore use supporting tools to distinguish between actions done with or without AI, or any other tools. We also added the emotions dimension, informed by prior literature (Cheng et al., 2023; Girardi et al., 2020; Graziotin et al., 2015; Madampe et al., 2025) and our pilot study (Section 2.7) to account for affective states that can reflect how emotions and cognitive programming behavior interplay (Girardi et al., 2020) and together shape the multidimensionality of developer-AI interactions. This results in our S–IASE model (State, Intention, Action, Supporting Tool, Emotion).
Stage 3: Dimension labeling. To develop the labels within each dimension, we drafted an initial set of labels based on prior studies (Mozannar et al., 2024; Amoozadeh et al., 2024; Russell, 1980; Gardella et al., 2024). Then, we allowed participants to add their custom labels during annotation if the predefined options were insufficient.
@@ 2.3. User Study Procedure
S1: Pre-survey. Each participant completed a pre-survey to obtain their consent to participate, demographic information (e.g., educational background), programming experience (especially on their AI use), and their prior knowledge on the overall topic of our designed tasks (e.g., familiarity with writing and executing SQL queries). We then used the information as a guide to assign participants to different groups (described in Section 2.5).
@@
S2: Activity. We provided two options for Setup according to participants’ preferences: (1) installing essential tools needed for the activity on their computers locally, such as task-specific programming language interpreters, a necessary database, a screen recording software, and our activity annotation tool. (2) installing a virtual machine that contains all these essential tools. Participants assigned to the AI-assisted group were explicitly informed that they could freely install and use their preferred AI assistants (e.g., ChatGPT or GitHub Copilot) without restriction, whereas participants in the non-AI group were instructed not to use any AI assistants during the tasks. Then, we introduced participants to the tasks and group assignment (i.e., AI or non-AI). Participants started screen recording and then performed their assigned task (Task Execution). After the task was completed, each participant watched their screen recording and used our custom annotation tool to break down the whole video into a sequence of intervals. For each interval, participants were asked to label their intentions, actions, supporting tools, and emotions.
S3: Post-survey. We first asked participants to upload their (1) screen recording of the task execution, (2) annotation results, and (3) implemented code. Then, we asked participants to reflect on their experience in completing the task in terms of cognitive load, self-efficacy, and challenges during task execution. Lastly, we collected their willingness to participate in the follow-up interview.
S4: Interview (optional). We invited interested participants to schedule an interview. The interviews helped triangulate our findings from the prior three stages. This additional data collection helped to develop a deeper understanding of participant behavior.
@@ 2.4. Activity Tasks
$\bullet$ JavaT: Java Task with MySQL Database Operations. This task involves setting up a MySQL database, implementing Java database operations to interact with the MySQL database, and solving several business queries over 4 subtasks.
@@ 2.5. Participants and Task Assignment
We recruited participants for our user study in two ways. The first was an in-class mode, where we gave a presentation in both undergraduate- and graduate-level software engineering courses to explain the study’s objectives and procedures. The second was recruiting participants through the research team’s internal connections and snowball sampling. To become a participant, one had to be over 18 years old and have prior programming experience. Our recruitment resulted in 90 participants, with 11 participants recruited via snowballing and 79 from the classroom. As compensation, 8 randomly selected participants received \$25 Amazon gift cards.
@@
We used stratified sampling to assign each participant to a group based on the following criteria: (1) If a participant had never or rarely used AI for programming, we assigned them to non-AI groups, i.e., nonAI-PyT and nonAI-JavaT. (2) Following the practice of Mozannar et al. (Mozannar et al., 2024), we assigned participants to a task according to their prior task-related knowledge in the pre-survey. We therefore created four participant groups: AI-PyT: participants in this group solved PyT and were allowed to use AI, nonAI-PyT: participants in this group solved PyT and were not allowed to use AI, AI-JavaT: participants in this group solved JavaT and were allowed to use AI, nonAI-JavaT: participants in this group solved JavaT and were not allowed to use AI. For data filtering, we did not consider a participant’s data if the screen recording did not capture the full screen or the annotation log file went missing. Moreover, to prevent the use of secondary devices (e.g., participants in non-AI groups using personal phones to access AI), researchers were present during all sessions. We also excluded four participants who did not complete the tasks during class. Researchers thoroughly examined the videos that participants recorded, and we excluded one participant in the non-AI groups who used AI in experiments. After filtering, we were left with 76 participants who completed all the study steps and complied with our instructions. Over 76% of participants reported 3 or more years of programming experience, with a median of over 4 years, and 75% reported prior internship or industry experience. Full demographic information is available in our replication package (Wu et al., 2026).